OpenStudy (anonymous):

if a car is traveling 70.0 mi/hr how many seconds will it take to go 35 km?

OpenStudy (anonymous):

the car is travelling by v=70.0mi/hr which equal (74.2km/hr) since v=d/t therefore t=d/v...........so,,,,,,,,,,,,,,,,,,,,t=35/74.2=0.47sec (Approx.)